# Unpacking Ambiguity: The Interaction of Polysemous Discourse Markers and Non-DM Signals

### v1

@@ Abstract
Jingni Wu Georgetown University jw2175@georgetown.edu Amir Zeldes Georgetown University az364@georgetown.edu
@@ 1 Introduction
Identifying and understanding discourse relations is fundamental to discourse comprehension. Discourse markers (DM) such as ‘and’, ‘because’, and ‘however’ have been widely recognized as the most typical indicator of coherence relations and are also referred to as discourse connectives or cue phrases Forbes-Riley et al. (2006). Early research focused on DMs as the sole device indicating relations, and their presence is often used to distinguish explicit and implicit relations Webber and Joshi (1998); Robaldo et al. (2008). However, more recent studies have shown that DMs account for only a small fraction of discourse relations, which can be signaled by reference (e.g. anaphora to indicate elaboration111Here and below we will assume discourse relation labels commonly used in Rhetorical Structure Theory Mann and Thompson (1988).), semantic (antonymy to indicate contrast), lexical (‘the next day’ can indicate temporal sequence like the DM ‘then’), morphological (past followed by present tense can also indicate sequence) and graphical cues (e.g. a question mark signaling a question relation).222For a full list of the non-DM signal types used in this study, see Appendix B. Such non-DM signals can be crucial for disambiguating otherwise ambiguous DMs, such as ‘since’, which can signal both cause and temporal circumstance relations. Taken together, DMs and such similar non-DM devices are referred to collectively as discourse relation signals Das and Taboada (2018a, b); Zeldes et al. (2025).
1.
2.
3.
2 Related Work
Although this pattern is very common in academic writing, little attention has been paid to the ways in which ambiguous DMs such as ‘while’ (which can also mean ‘during a time that…’) resolve to a unique intepretation thanks to co-occurring signals in this manner, and the joint use of DMs and signals remains a complex question.
Non-DM signals can 1) overlap with DMs in meaning, potentially leading to redundancy, 2) co-occur with DMs but function independently (potentially signaling multiple distinct relations), and 3) may complement DMs in specific types of relations and environments Hoek et al. (2018). Recent studies have begun to explore the underlying triggers of the ‘DM + other signals’ phenomenon. Das and Taboada (2019) suggested that such combinations may arise from the inherent ambiguity of certain DMs which can signal various relations. For example, the DM and can mark additive list and temporal sequence relations, among other options, as illustrated in the following examples from GUM:
Building on this, researchers have introduced the concept of ‘marking strength’ or ‘signaling strength’ of DMs, which can be assessed by the number and frequency of discourse relations they can signal Asr and Demberg (2012). Zeldes and Liu (2020) proposed the ‘delta-softmax’ metric, which quantifies prediction accuracy degradation for a trained neural model when a word is removed to estimate its signaling strength for a relation, providing empirical validation of an intuitive graded ‘signaly-ness’ phenomenon. For instance, ‘but’ could be significantly less ambiguous than ‘and’ as a DM, in that removing ‘but’ would make the relation much harder to predict than removing ‘and’.
In addition, combinations of DMs and non-DM signals vary across relation types, but they are not necessarily driven by inherent semantics Das and Taboada (2019). That is to say, certain relations tend to prefer either DM-only or DM-plus-signal combinations, which is partly influenced by the inherent semantics of the discourse relations themselves (e.g., weakly connected sentences), but also appears to reflect an independent pragmatic strategy for ensuring clarity of the writer’s intention.
3 Data
This study uses the Georgetown University Multilayer (GUM) Corpus which consists of 16 spoken and written, informal and formal style English text types Zeldes (2017) (see corpus details in Appendix A). The corpus originally contained RST annotations, which were recently extended based on Enhanced Rhetorical Structure Theory (eRST, Zeldes et al. 2025), adding annotated DMs and seven types of non-DM signals based on the taxonomy proposed by Das and Taboada Das and Taboada (2018a), as well as adding multiple concurrent and tree breaking relations edges to the initial RST trees. With over 250,000 tokens, this is currently the largest dataset annotated for DMs and non-DM signals of discourse relations.
4 Polysemy of DMs
The ambiguous nature of DMs arises from their one-to-many relationship with discourse relations. DMs that can signal multiple relations, such as ‘and’, are often described as ‘weak signals’ Asr and Demberg (2012); Das and Taboada (2019); Crible (2020), as they do not map consistently to a single meaning, in contrast to unambiguous DMs such as ‘despite’, which always marks a concession.
5 DM-Signal Co-occurrences
5.1 General Distribution
Across 16 genres, 21,435 discourse relations are annotated in our data, of which 1,372 (6.4%) are indicated by both DMs and non-DM signals. This result aligns fairly closely with Das and Taboada (2018a)’s finding for Wall Street Journal news (7.55%). However as suspected, we observe substantial variation across genres (see Figure 1): essay (8.7%), bio (8.6%), and whow (8.5%) show a higher proportion of DM-signal co-occurrence, whereas conversation has the lowest proportion (3.9%).
| Total counts | 1092 | 229 | 42 | 6 | 1 | 1 | 1 |
| Proportion | 79.55% | 16.7% | 3.1% | 0.44% | 0.07% | 0.07% | 0.07% |
. by using a second order Rao and Scott (1981) … correction
The top three most frequently used signal types in co-occurrences are ‘semantic’, ‘syntactic’, and ‘lexical’ across genres, though different genres favor different types of signals, in part due to the format of texts. In spoken genres such as vlog, conversation, and court, there is a large amount of ‘reference’ signals used along with DMs to indicate relations such as elaboration (see Figure 2). Trivially, ‘graphical’ signals such as quotation marks to signal attribution cannot occur in spoken language and are restricted to written data.
5.2 Polysemous DMs and Signal Patterns
The DM ‘so’ has the highest polysemy score across all genres in our dataset, while the DM ‘for’ exhibits the most diverse range of accompanying signals (see Table 2). Here, ‘diversity’ 333Since DM frequency varies across genres, we normalized diversity by dividing the number of unique co-occurring signal types by the square root of total DM occurrences. This accounts for diminishing returns and prevents frequent DMs from being unfairly penalized. refers to the number of distinct non-DM signal types that co-occur with a given DM, including individual signal types (e.g. ‘semantic’) and combinations of multiple types (e.g. ‘semantic + lexical’).
| DM | non-DM signal diversity |
| for | 29.50 |
| and | 26.64 |
| if | 25.00 |
| by | 20.80 |
| when | 19.00 |
Looking at patterns rather than counts of signals in more detail, certain signal types consistently co-occur with highly polysemous DMs, suggesting that these signals play a crucial role in disambiguating them. For example, ‘lexical’ and ‘syntactic’ signals frequently appear across multiple cases and are more likely to be combined with other signal types, reinforcing their role in guiding interpretation (see Table 3).
| DM | Top 3 co-occurring Types | Top 3 most frequent combinations |
| as | syntactic, lexical, morphological | (lexical + semantic) |
5.3 Signal Combinations and Genre Effects
The non-DM signals that co-occur with polysemous DMs exhibit diverse combination patterns, which vary across genres. A single DM may be more likely to be paired with entirely different signals depending on the genre. For example, the DM ‘and’ is most frequently used with ‘lexical_chain’ signals (a subtype of ‘semantic’) signals, see example 5.3) in nearly all genres, except for vlog, bio, whow, conversation, and podcast (see Appendix Table 7). Here the lexical relation between the related items ‘information’ and ‘content’ forms a semantic signal next to ‘and’ to indicate that the two clauses are part of a list.
This variation can be attributed to the nature of spoken genres such as vlogs and conversations, which emphasize audience interaction and shared common ground. In these contexts, indicative words and personal references are more commonly used to enhance engagement and coherence. Similarly, other spoken genres tend to favor ‘reference’ signals, particularly ‘personal references’, using chains of pronouns to help the audience recall previously mentioned content. Semantic signals in the genre podcast show a particularly strong use of meronymy, using words in a part-whole relationship alongside the polysemous ‘and’ to indicate elaborations on complex information.
In addition, ‘and’ tends to use combined signals more frequently than other polysemous DMs, dovetailing with our initial hypothesis about non-DMs compensating for ambiguous DMs. Notably, in almost all genres where ’and’ is the most polysemous DM, it co-occurs with multiple signals, except for the genre essay, where it primarily appears by itself or with a single signal type. Among all signal combinations, the most frequent combined signal set for ‘and’ is ‘reference’ + ‘semantic’, i.e. anaphora and lexical relations between words in the units joined by ‘and’. Interestingly, letter is the only genre where the most polysemous DM is ‘as’, yet it does not co-occur with any additional non-DM signals. Looking at its instances, nearly 65% are used to indicate mode relations (manner/means), as opposed to only 32.2% in the rest of the corpus, suggesting that this usage may simply be more predictable as a default in letters – the most common sense in the remaining genres is indicating a temporal circumstance, similarly to ‘when’.
Among the genres, academic, reddit, and court seem to have larger variance, indicating that DMs used in these genres experience the most significant shifts in polysemy compared to their global usage. These genres may have DMs that behave very differently in terms of polysemy compared to their global usage. In contrast, DMs in fiction, podcast, and letter appear to behave similarly locally and globally.
6 Discussion
This study investigates the relationship between DM polysemy, the number and diversity of co-occurring non-DM signals, and the role of genre in these interactions. Our findings partially support the hypothesis that polysemous DMs exhibit more diverse non-DM signal patterns but do not necessarily co-occur with a greater number of non-DM signals. Moreover, genre greatly shapes DM polysemy, with significant variations in DM entropy and signal usage.
The regression analysis suggests that polysemous DMs are more likely to co-occur with more diverse non-DM signals. However, this relationship varies by genre, with spoken genres (e.g. court, podcast, vlog) showing a stronger dependence on non-DM signals to disambiguate polysemous DMs. This may reflect the unique cognitive and interactional constraints of speech. Spoken discourse requires real-time processing, and speakers often experience pressure to maintain fluency while managing limited cognitive resources for lexical retrieval Clark (2002). As a result, they may rely more on non-DM signals to support discourse coherence, even when these are partially redundant. Furthermore, DMs in spoken texts often serve interactive and procedural functions beyond marking semantic relations, such as managing turn-taking or structuring discourse Clark and Tree (2002). This functional difference may explain why speech genres exhibit stronger correlations between DM polysemy and signal use (numbers and types), as speakers may compensate for ambiguity and uncertainty in perception through additional diverse discourse cues.
On the other hand, some genres exhibit little to no significant correlation among DM polysemy, the number and the diversity of non-DM signals, suggesting that different discourse contexts may impose different constraints on how DMs interact with non-DM signals. Additionally, we identified DMs whose polysemy scores are highly shifted across genres, such as, ‘so’, ‘in’, ‘with’, ‘given’, ‘indeed’, and ‘while’. This finding suggests that certain polysemous DMs are more sensitive to contextual variation, whereas others maintain stable meanings across different discourse settings.
This study does not fully account for the distribution of different discourse relations, which can further shape the observed patterns of polysemy and signal co-occurrence. Prior research has demonstrated that certain non-DMs are more commonly used to disambiguate DMs in specific relations, such as contrast and consequence Crible and Demberg (2020), and different relations may vary in their sensitivity to signals, with some relations being more reliant on co-occurring non-DM cues for disambiguation. Moreover, the compatibility between DMs and specific signals may play a greater role in guiding interpretation than sheer signal frequency. Future research could explore how different discourse relations condition the use of non-DM signals with polysemous DMs, providing a more fine-grained understanding of signal-based disambiguation mechanisms.
Appendix A GUM Information
| Text type | Source | Docs | Tokens |
| Academic writing | Various | 18 | 17,169 |
| Biographies | Wikipedia | 20 | 18,213 |
| CC Vlogs | YouTube | 15 | 16,864 |
| Conversations | UCSB Corpus | 15 | 17,932 |
| Courtroom transcripts | Various | 9 | 11,148 |
| Essays | Various | 9 | 10,842 |
| Fiction | Various | 19 | 17,511 |
| Forum | 18 | 16,364 | |
| How–to guides | wikiHow | 19 | 17,081 |
| Interviews | Wikinews | 19 | 18,196 |
| Letters | Various | 12 | 9,989 |
| News stories | Wikinews | 24 | 17,186 |
| Podcasts | Various | 10 | 11,986 |
| Political speeches | Various | 15 | 16,720 |
| Textbooks | OpenStax | 15 | 16,693 |
| Travel guides | Wikivoyage | 18 | 16,515 |
| Total GUM | 255 | 250,409 |
Appendix B Non-DM signal types
In this paper we follow the taxonomy of non-DM signal types proposed by Zeldes et al. (2025), which distinguises eight major classes with a total of 45 subtypes, illustrated in Table 6.

### v2

@@ Abstract
Jingni Wu Georgetown University jw2175@georgetown.edu Amir Zeldes Georgetown University amir.zeldes@georgetown.edu
@@ 1 Introduction
Identifying and understanding discourse relations is fundamental to discourse comprehension. Discourse markers (DM) such as ‘and’, ‘because’, and ‘however’ have been widely recognized as the most typical indicator of coherence relations and are also referred to as discourse connectives or cue phrases Forbes-Riley et al. (2006). Early research focused on DMs as the sole device indicating relations, and their presence is often used to distinguish explicit and implicit relations Webber and Joshi (1998); Robaldo et al. (2008). In applied Natural Language Processing (NLP) they also remain the focus of research on automatic detection of discourse relation signaling, as evidenced by the series of DISRPT (Discourse Relation Parsing and Treebanking, see Braud et al. 2024) shared tasks including DM detection as a track. Since such markers come from a closed list, systems can target only these words or phrases Yu et al. (2019), then focus on disambiguation, with recent system scores achieving over 93% F1-scores for English Liu et al. (2023).
@@
However, more recent studies have shown that DMs account for only a small fraction of discourse relations, which can be signaled by reference (e.g. anaphora to indicate elaboration111Here and below we will assume discourse relation labels commonly used in Rhetorical Structure Theory Mann and Thompson (1988). Our definition of what constitutes anaphoric reference aligns with Zeldes (2022).), semantic (antonymy to indicate contrast), lexical (‘the next day’ can indicate temporal sequence like the DM ‘then’), morphological (past followed by present tense can also indicate sequence) and graphical cues (e.g. a question mark signaling a question relation). In this paper we follow the taxonomy of non-DM signal types proposed by Zeldes et al. (2025), which distinguishes eight major classes with a total of 45 subtypes, illustrated in Table 1. Such non-DM signals can be crucial for disambiguating otherwise ambiguous DMs, such as ‘since’, which can signal both cause and temporal circumstance relations. Taken together, DMs and such similar non-DM devices are referred to collectively as discourse relation signals Das and Taboada (2018a, b); Zeldes et al. (2025).
@@
1.
2.
3.
2 Related Work
Although this pattern is very common in academic writing, little attention has been paid to the ways in which ambiguous DMs such as ‘while’ (which can also mean during a time that…) resolve to a unique intepretation thanks to co-occurring signals in this manner, and the joint use of DMs and signals remains a complex question.
Non-DM signals can 1) overlap with DMs in meaning, potentially leading to redundancy, 2) co-occur with DMs but function independently (potentially signaling multiple distinct relations), and 3) may complement DMs in specific types of relations and environments Hoek et al. (2018). Recent studies have begun to explore the underlying triggers of the DM + other signals phenomenon. Das and Taboada (2019) suggested that such combinations may arise from the inherent ambiguity of certain DMs which can signal various relations. For example, the DM and can mark additive list and temporal sequence relations, among other options, as illustrated in the following examples from GUM:
Building on this, researchers have introduced the concept of marking strength or signaling strength of DMs, which can be assessed by the number and frequency of discourse relations they can signal Asr and Demberg (2012). Zeldes and Liu (2020) proposed the delta-softmax metric, which quantifies prediction accuracy degradation for a trained neural model when a word is removed to estimate its signaling strength for a relation, providing empirical validation of an intuitive graded signaly-ness phenomenon. For instance, ‘but’ could be significantly less ambiguous than ‘and’ as a DM, in that removing ‘but’ would make the relation much harder to predict than removing ‘and’.
In addition, combinations of DMs and non-DM signals vary across relation types, but they are not necessarily driven by inherent semantics Das and Taboada (2019). That is to say, certain relations tend to prefer either DM-only or DM-plus-signal combinations. This is partly influenced by the inherent semantics of the discourse relations themselves (e.g., weakly connected sentences), but also appears to reflect an independent pragmatic strategy for ensuring clarity of the writer’s intention.
3 Data
This study uses the Georgetown University Multilayer (GUM) Corpus which consists of 16 spoken and written, informal and formal style English text types Zeldes (2017) (see corpus details in Appendix A). The corpus originally contained Rhetorical Structure Theory (RST) annotations, which were recently extended based on Enhanced Rhetorical Structure Theory (eRST, Zeldes et al. 2025), adding annotated DMs and seven types of non-DM signals based on the taxonomy proposed by Das and Taboada (2018a), as well as adding multiple concurrent and tree breaking relations edges to the initial RST trees. With over 250,000 tokens, this is currently the largest dataset annotated for DMs and non-DM signals of discourse relations.
Since this study relies on accurate discourse annotations, we also report how quality was assured. Inter-annotator agreement studies on GUM showed F1 scores of 92.3 for DM identification and 90 for relation association (36 docs, 32K tokens). For non-DM signals, many types were automatically derived from gold syntax and coreference annotations, with others manually corrected or added. On a subset of documents, human–human agreement yielded an F1 of about 0.80.
4 Polysemy of DMs
The ambiguous nature of DMs arises from their one-to-many relationship with discourse relations. DMs that can signal multiple relations, such as ‘and’, are often described as weak signals Asr and Demberg (2012); Das and Taboada (2019); Crible (2020), as they do not map consistently to a single meaning, in contrast to unambiguous DMs such as ‘despite’, which always marks a concession.
5 DM-Signal Co-occurrences
5.1 General Distribution
Across 16 genres, 21,435 discourse relations are annotated in our data, of which 1,372 (6.4%) are indicated by both DMs and non-DM signals. This result aligns fairly closely with Das and Taboada (2018a)’s finding for Wall Street Journal news (7.55%, see also Liu and Zeldes 2019). However as suspected, we observe substantial variation across genres (see Figure 1): essay (8.7%), bio (8.6%), and whow (how-to guides form Wikihow, 8.5%) show a higher proportion of DM-signal co-occurrence, whereas conversation has the lowest proportion (3.9%).
| Total counts | 1092 | 229 | 42 | 6 | 1 | 1 | 1 |
| Proportion | 79.55% | 16.7% | 3.1% | 0.44% | 0.07% | 0.07% | 0.07% |
. by using a second order Rao and Scott (1981) … correction
The top three most frequently used signal types in co-occurrences are semantic, syntactic, and lexical across genres, though different genres favor different types of signals, in part due to the format of texts. In spoken genres such as vlog, conversation, and court, there is a large amount of reference signals used along with DMs to indicate relations such as elaboration (see Figure 2). Trivially, graphical signals such as quotation marks to signal attribution cannot occur in spoken language and are restricted to written data.
5.2 Polysemous DMs and Signal Patterns
The DM ‘so’ has the highest polysemy score across all genres in our dataset, while the DM ‘for’ exhibits the most diverse range of accompanying signals (see Table 3). Here, diversity 222Since DM frequency varies across genres, we normalized diversity by dividing the number of unique co-occurring signal types by the square root of total DM occurrences. This accounts for diminishing returns and prevents frequent DMs from being unfairly penalized. refers to the number of distinct non-DM signal types that co-occur with a given DM, including individual signal types (e.g. semantic) and combinations of multiple types (e.g. semantic + lexical).
| DM | non-DM signal diversity |
| for | 29.50 |
| and | 26.64 |
| if | 25.00 |
| by | 20.80 |
| when | 19.00 |
Looking at patterns rather than counts of signals in more detail, certain signal types consistently co-occur with highly polysemous DMs, suggesting that these signals play a crucial role in disambiguating them. For example, lexical and syntactic signals frequently appear across multiple cases and are more likely to be combined with other signal types, reinforcing their role in guiding interpretation (see Table 4).
| DM | Top 3 co-occurring Types | Top 3 most frequent combinations |
| as | syntactic, lexical, morphological | (lexical + semantic) |
5.3 Signal Combinations and Genre Effects
The non-DM signals that co-occur with polysemous DMs exhibit diverse combination patterns, which vary across genres. A single DM may be more likely to be paired with entirely different signals depending on the genre. For example, the DM ‘and’ is most frequently used with lexical_chain signals (a subtype of semantic) signals, see example 5.3) in nearly all genres, except for vlog, bio, whow, conversation, and podcast (see Appendix Table 7). Here the lexical relation between the related items ‘information’ and ‘content’ forms a semantic signal next to ‘and’ to indicate that the two clauses are part of a list.
This variation can be attributed to the nature of spoken genres such as vlogs and conversations, which emphasize audience interaction and shared common ground. In these contexts, indicative words and personal references are more commonly used to enhance engagement and coherence. Similarly, other spoken genres tend to favor reference signals, particularly personal references, using chains of pronouns to help the audience recall previously mentioned content. Semantic signals in the genre podcast show a particularly strong use of meronymy, using words in a part-whole relationship alongside the polysemous ‘and’ to indicate elaborations on complex information.
In addition, ‘and’ tends to use combined signals more frequently than other polysemous DMs, dovetailing with our initial hypothesis about non-DMs compensating for ambiguous DMs. Notably, in almost all genres where ‘and’ is the most polysemous DM, it co-occurs with multiple signals, except for the genre essay, where it primarily appears by itself or with a single signal type. Among all signal combinations, the most frequent combined signal set for ‘and’ is reference + semantic, i.e. anaphora and lexical relations between words in the units joined by ‘and’. Interestingly, letter is the only genre where the most polysemous DM is ‘as’, yet it does not co-occur with any additional non-DM signals. Looking at its instances, nearly 65% are used to indicate mode relations (manner/means), as opposed to only 32.2% in the rest of the corpus, suggesting that this usage may simply be more predictable as a default in letters – the most common sense in the remaining genres is indicating a temporal circumstance, similarly to ‘when’.
Among the genres, academic, reddit, and court seem to have larger variance, indicating that DMs used in these genres experience the most sizable shifts in polysemy compared to their global usage. These genres may have DMs that behave very differently in terms of polysemy compared to their global usage. In contrast, DMs in fiction, podcast, and letter appear to behave similarly locally and globally.
6 Conclusion
This study investigates the relationship between DM polysemy, the number and diversity of co-occurring non-DM signals, and the role of genre in these interactions. Our findings partially support the hypothesis that polysemous DMs exhibit more diverse non-DM signal patterns but do not necessarily co-occur with a greater number of non-DM signals. Moreover, genre greatly shapes DM polysemy, with significant variations in DM entropy and signal usage. Spoken genres (e.g. court, podcast, vlog) show a stronger dependence on non-DM signals to disambiguate polysemous DMs, while some written genres exhibit little or no correlation. This pattern likely reflects cognitive and interactional pressures in speech, where speakers must maintain fluency under real-time constraints Clark (2002) and often deploy additional cues to support coherence. Moreover, DMs in spoken discourse frequently serve interactive functions, such as managing turn-taking or structuring talk Clark and Tree (2002), which may further increase their co-occurrence with diverse signals.
These findings challenge theoretical views of DMs in frameworks that assume that DM marking means we do not need to consider other types of signals, such as in the Penn Discourse Treebank framework, where alternative lexicalizations marking a relation are generally only considered if a DM is absent Prasad et al. (2018). They also suggest a consequence for treating DMs and other types of markers as categorically consistent across different types of text: in practice, we find great variation in the extent and types of signaling present based on genre.
On the other hand, some genres exhibit little to no significant correlation among DM polysemy, the number and the diversity of non-DM signals, suggesting that different discourse contexts may impose different constraints on how DMs interact with non-DM signals. Additionally, we identified DMs whose polysemy scores are highly shifted across genres, such as, ‘so’, ‘in’, ‘with’, ‘given’, ‘indeed’, and ‘while’. This finding suggests that certain polysemous DMs are more sensitive to contextual variation, whereas others maintain stable meanings across different discourse settings. Further research is needed to understand the extent to which the picture of genre variation presented here is comprehensive, which could be carried out with new eRST data on unusual genres that has recently become available (for example in the GENTLE corpus, Aoyama et al. 2023, which includes annotations for poetry, legal writing, and more).
7 Discussion
This study does not fully account for the distribution of different discourse relations, which can further shape the observed patterns of polysemy and signal co-occurrence. Prior research has demonstrated that certain non-DMs are more commonly used to disambiguate DMs in specific relations, such as contrast and consequence Crible and Demberg (2020), and different relations may vary in their sensitivity to signals, with some relations being more reliant on co-occurring non-DM cues for disambiguation. Moreover, the compatibility between DMs and specific signals may play a greater role in guiding interpretation than sheer signal frequency. Future work should therefore examine how relation type conditions the use of non-DM signals with polysemous DMs, and expand analysis to larger silver-standard multilayer corpora such as AMALGUM (A Machine Annotated Lookalike of GUM, Gessler et al. 2020), enriched with automatic annotation of discourse relations and signals, which would be less accurate, but mitigate the problem of data sparseness.
Beyond theoretical implications, these findings also have practical relevance for NLP. Current discourse parsers often treat explicit relations with DMs as straightforward, yet our results show that polysemous markers frequently rely on co-occurring signals. Incorporating such cues could improve discourse relation classification and domain adaptation, while also enhancing explainability in downstream tasks such as summarization or dialogue systems.
Appendix A GUM Information
| Text type | Source | Docs | Tokens |
| Academic writing | Various | 18 | 17,169 |
| Biographies | Wikipedia | 20 | 18,213 |
| CC Vlogs | YouTube | 15 | 16,864 |
| Conversations | UCSB Corpus | 15 | 17,932 |
| Courtroom transcripts | Various | 9 | 11,148 |
| Essays | Various | 9 | 10,842 |
| Fiction | Various | 19 | 17,511 |
| Forum | 18 | 16,364 | |
| How–to guides | wikiHow | 19 | 17,081 |
| Interviews | Wikinews | 19 | 18,196 |
| Letters | Various | 12 | 9,989 |
| News stories | Wikinews | 24 | 17,186 |
| Podcasts | Various | 10 | 11,986 |
| Political speeches | Various | 15 | 16,720 |
| Textbooks | OpenStax | 15 | 16,693 |
| Travel guides | Wikivoyage | 18 | 16,515 |
| Total GUM | 255 | 250,409 |
Appendix B Signal Patterns for "and" by Genre